OpenStudy (anonymous):

If 12.03 kJ of heat are required to melt 36.0 grams of water, what is the heat of fusion of water? 2.99 kJ/mol 6.02 kJ/mol 12.04 kJ/mol 60.20 kJ/mol @Somy @Abhisar

OpenStudy (abhisar):

Hello !

OpenStudy (abhisar):

Can u calculate the moles of water in 36 grams ?

OpenStudy (anonymous):

no :( am bad at chem @Abhisar

OpenStudy (abhisar):

Do u know the molar mass of water ?

OpenStudy (anonymous):

18.02 @Abhisar

OpenStudy (abhisar):

good !

OpenStudy (anonymous):

@somy now am confused again

OpenStudy (abhisar):

See it's simple, 12.03 kJ/mol required for 36 grams or 2 moles

OpenStudy (abhisar):

Then for one mole how much heat is required ?

OpenStudy (abhisar):

Define latent heat of fusion

OpenStudy (anonymous):

so do i divide 12.03 by 36.0 @Somy @Abhisar

OpenStudy (anonymous):

but it doesnt give me the answer i need

OpenStudy (abhisar):

See the unit @Somy kJ/m = kJ per mole

OpenStudy (somy):

mass= is in Kg so change it to Kg

OpenStudy (somy):

so = 36/1000= 0.036 kg now u can use the formula

OpenStudy (anonymous):

i got 0.299

OpenStudy (somy):

probably @Abhisar im not having calculator atm

OpenStudy (somy):

lol i think i confused everyone +_+ lets start it over

OpenStudy (anonymous):

@somy i got the answer wrong :(

OpenStudy (abhisar):

I think it should be B

OpenStudy (somy):

how did u come up with it?

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar Then for one mole how much heat is required ? \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar See it's simple, 12.03 kJ/mol required for 36 grams or 2 moles \(\color{blue}{\text{End of Quote}}\)

OpenStudy (somy):

1 mole of H2O= 18

OpenStudy (somy):

12.03/2 =

OpenStudy (somy):

wait lol

OpenStudy (somy):

@Abhisar now im confused isn't Q= 12.03 ?

OpenStudy (somy):

but they are asking for Latent heat of fusion

OpenStudy (somy):

wait a sec are they just indirectly asking for Q and by bringing up 'fusion' they are confusing us? @Abhisar

OpenStudy (somy):

@veronica.montanez are you taking Cambridge system?

OpenStudy (anonymous):

@Somy no

OpenStudy (somy):

hmmm @Abhisar take over this - i don't really get what they are asking for coz for Latent heat of Fusion =number is real big

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar See it's simple, 12.03 kJ/mol required for 36 grams or 2 moles \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar Then for one mole how much heat is required ? \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abmon98):

eat of fusion is the amount of heat energy required to change the state of a substance from solid to liquid. Q=m*Delta f Delta f unit is KJ/mol 36 grams is 2 moles of water. Delta F= 12.04/2=6.02KJ per mole